if the kinetic energy of a 6.2 kg object is 36.5 J what is the objects velocity


1
Expert's answer
2021-05-13T14:26:22-0400

The equation for kinetic energy is


"E=\frac 12 mv^2."

The velocity, therefore, is


"v=\sqrt{\frac{2E}{m}}=3.43\text{ m/s}."